1(a) a vessel containing an unknown volume of argon gas in cm3 at temperature 300f and pressure 101kpa was inverted and placed in cold ethanol. As the gas contracted, the pressure was quadrupled while the temperature reduced by one-third. What is the final volume of gas in dm3?
(B.) Calculate the root mean square of bromine molecule at temperature 23 degree Celsius and at pressure 1.00 atm. What is the root mean square of the same gas at temperature 23 degree Celsius and pressure 1.5 atm
Expert's answer
1
Expert's answer
2019-04-30T04:30:48-0400

Solution:

a)pV=nRT

nR is unchanged; nR=1; T1=422 K; T2=300/3=100 Fahrenheit=310,9 K.

V1=1*422/101000=0,0042 L

V2=1*310,9/404000=0,00077 L

Answer: 0,00077 L

b)Urms="\sqrt{\smash[b]{3p/\rho}}"

1.​ "\rho" =p*M/R*T; M(Br2)=2*80=160 grams/mol; R=0,08206 atm*L/mol*K; T=23+273=296 K

ρ=1*160/0,08206*296=6,587 grams/L;

Urms="\sqrt{\smash[b]{3*1/6,587}}" =0,4554 m/s

2.ρ=1,5*160/0,08206*296=9,88 grams/L;

Urms="\sqrt{\smash[b]{3*1,5/9,88}}" =0,4555 m/s.

Answer: 0,4554 m/s; 0,4555 m/s.